Please show work.
Using reasonable estimates of height differences, calculate the blood pressure difference between each of the following and the heart (specifically, pressure at the body part minus pressure at the heart): head, hand, foot. Write the answers in Pa and in mm Hg. Assuming that the pressure at the heart is 120 mm Hg, calculate the pressure at the head, at the hand, and at the foot.
Answer #1

Assuming the height of person=180cm

heart is situated at 60cm from head

foot is at 120cm from heart

Pressure at heart =120mm of Hg

Let air density=1.225kg/m^3

Pressure at head will be lesser by .6*1.225/13.6

=0.054m of Hg=54mm

Pressure at foot will be more by 1.2*1.225/13.6

=.108m of Hg=108mm

Pheart-Phead=54mm of Hg=10^5*54/760=7105Pa

Pheart-Pfoot=108mm of Hg=10^5*108/760=-14210Pa